# Going Proactive and Explanatory Against Malware Concept Drift

### v1

@@ Abstract
1 Introduction
@@
Existing methods for drift detection, however, fall short in two main aspects. Firstly, they often falsely neglect the patterns which the targeted classifier depends on. For example, the CADE detector [75] leverages an independent autoencoder to learn a distance function, identifying drifts by the distances with training data. Such misalignment can lead to inefficiency in detecting model-specific drifts, especially when dealing with complex classifier feature spaces, as validated by our experiments in Section 6.2. Secondly, a constant involvement of training data during the testing phase introduces practical challenges. For instance, methods use training points to query reference uncertainties [7, 33] or search for nearest neighbors [14]. Such practices are problematic particularly in local deployment scenarios, where managing large training datasets raises storage and security concerns [53].
@@
We develop a system, called Dream, distinguished by concept-based drift detection and explanatory adaption for malware classification. To accurately learn the behavioral concepts, we combine supervised concept learning [31] with unsupervised contrastive learning [13]. Our drift detector is thus a semi-supervised model structured with an autoencoder. Within this model, elements in the latent space are representative of behavioral concepts that can be labeled, such as remote control and stealthy download. Specifically, Dream achieves model sensitive and training data independent drift detection with the concept reliability measure. This unique measure is based on the intuition that if a sample is not subject to drift, then the concepts embedded in our detector should be reliable; consistency should be reflected in similar classifier predictions for the original and reconstructed samples.
@@
We evaluate Dream with two distinct malware datasets [75, 69] and three widely-used malware classifiers [4, 44, 47]. Evaluation results demonstrate that Dream outperforms three existing drift detection methods [7, 75, 14] in malware classification, achieving an average improvement of 11.5%, 12.0%, and 13.6% in terms of the AUC  [^1] The AUC specifically refers to the Area Under the Receiver Operating Characteristic (ROC) Curve [19] in this paper. metric. For drift adaptation, Dream notably enhances the common retraining-based approach, increasing the F1-score by 168.4%, 73.5%, 22.2%, 19.7%, and 6.0% when the human analysis budget is set at $10$, $20$, $30$, $40$, and $100$, respectively. A standout observation is our increased advantage at lower human labeling costs. For instance, analysts can analyze over 80% fewer samples while still achieving a test accuracy of $0.9$ for the updated classifier.
@@
2 Malware Concept Drift
In this section, we delve into the background of malware concept drift. Firstly, we discuss two types of drift for learning-based malware applications. Then, we outline the widely employed techniques for drift detection and drift adaptation (with supplements available in Section 9). Finally, we highlight key studies in this domain, underscoring their unique research orientations and inherent limitations.
2.1 Root Cause and Categorization
@@
2.2 Drift Detection
Drift detection is essential for stakeholders to determine when to update the classifier. Two primary Out-of-Distribution (OOD) detection methods [74] are chiefly employed.
1) Uncertainty Estimation: Uncertainty estimation gauges a model’s prediction confidence. Techniques like Bayesian Neural Networks [36] and Deep Ensembles [38] inherently quantify this measure, though they require architectural changes and can be computationally demanding. Standard DNNs often use heuristics like softmax-derived uncertainty, which can be misleading as overfit DNNs may assign high softmax values to unfamiliar data, falsely indicating confidence [28, 49].
2) Nonconformity Scoring: Nonconformity scoring, in contrast, evaluates how unusual new data is compared to a calibration set [33]. It involves a distance function and a predictor for representing calibration and test data. While traditional techniques like conformal prediction methods depend heavily on the calibration set, recent approaches use contrastive learning to refine distance functions [75, 14, 66]. However, their model independent nature can lead to inconsistent results across different architectures.
In conclusion, both uncertainty estimation and nonconformity scoring offer valuable insights, but their effectiveness varies depending on the model’s nature and the calibration data’s quality. A continuous evaluation, possibly integrating both methods, is crucial for comprehensive drift detection.
2.3 Drift Adaptation
Drift adaptation in malware classification traditionally involves manual feature engineering and periodic model retraining [79, 17, 8, 15]. While effective, these methods can be specific to certain features and require extensive manual labeling. Recently, several novel adaptation strategies have emerged from advancements in drift detection methodologies.
One such strategy, classification with rejection, adopts a conservative approach where ambiguous samples are temporarily withheld for expert analysis to minimize misclassification risks [33, 7]. Another strategy, inspired by active learning, introduces an analysis budget to prioritize a subset of drifting samples for human review and subsequent integration into retraining [14]. However, it’s notable that existing studies all focus on the adaptation for intra-class drift, where the number of output classes is unchanged. Additionally, there’s growing interest in drift explanation, where the focus is on understanding the reasons behind drift samples to streamline manual analysis [75, 25, 24]. Nevertheless, the post-analysis application of these explanations in updating classifiers remains underexplored.
While classification with rejection provides a safeguard against immediate threats, its efficacy in long-term resilience is limited. Current strategies that adopt active learning focus on intra-class drift, adapting to subtle malware evolutions. Yet, addressing inter-class drift and fully harnessing explanatory analysis for classifier updates still need more research.
2.4 Summary of Existing Work
Advancements in drift detection and adaptation techniques are crucial for combating malware drifts effectively. Focusing on these aspects, three notable studies have contributed significantly to the field. As outlined in Table 1, we discuss their methodologies and comparative research focuses below, and technical details of their detectors can be found in Section 3.3.
@@ 2.4 Summary of Existing Work
Table 1: Summary of existing work in terms of the detection method, adaptation method, and the applicable scope.
@@
3 Problem Setup
In this section, we define our research scope, formalize important components in the active learning-based strategy, and characterize current drift detectors with our formalization.
3.1 Research Scope
@@
Distinguishing our method from HCC, we emphasize model-agnostic techniques, which are adept at operating effectively without modifications to the underlying classifier’s architecture. Recognizing that intra-class drift scenarios have already been extensively explored, our work primarily addresses the challenges associated with inter-class drifts. That is, we aim to delve deeply into the detection and adaptation strategies for drifts caused by unseen malware families. Nevertheless, it’s important to note that our proposed solution is not exclusively tailored to addressing inter-class drifts. The generalizability of our approach will also undergo a detailed evaluation in Section 7.
@@ 3.2 Formalization
3.3 Characterizing Current Drift Detectors
@@ 4 Motivation and Challenge
Table 2: Intra-class drift detection performance of existing works. For the state-of-the-art method HCC, we also make a separate study of the two items in its drift scoring function.
4.1 Lessons from Current Detectors
@@
Evaluation Method. For malware detection tasks, we leverage malware samples from the Malradar dataset [69], which covers 4,410 malware across 180 families including singletons. Moreover, we make efforts to collect 43,641 benign samples in the same year period from 2015 to 2020. Our dataset has been carefully designed to mitigate spatial bias and temporal bias [50]. For the classifier, we use the Drebin features and an MLP model (detailed in Section 6.1). Regarding baselines, we take into account the vanilla probability-based detector (as depicted in Equation 4) and the three innovative detectors formalized in Section 3.3. During the evaluation phase, we assign positive drift labels to samples where predictions are incorrect, following the rationale that these misclassified samples are particularly valuable for adapting the classifier. This approach differs from inter-class drift detection, where positive drift labels are typically assigned based on the presence of unseen families, as utilized in existing work [75]. Then we compute the AUC metric using drifting scores and the ground truth drift labels.
@@
Key Findings. Our evaluation results are shown in Table 2, from which we have identified four key findings. 1) The performance order in our experiments, i.e., HCC, Transcendent, and Probability all outperforming CADE, highlights the essential role of model sensitivity in effective drift detection. 2) Although CADE is not specifically tailored for intra-class drift, its superior performance in a non-model sensitive comparison with the $\text{HCC}_{hc}$ indicates the potential advantages of data autonomy and its autoencoder-based structure. 3) HCC surpassing Transcendent in performance indicates that learning a distance metric, rather than relying on calibration, can be more effective and efficient in complex models like DNNs. This observation contrasts with Transcendent’s findings on SVM models and is further supported by the efficiency of the detector’s training process, which can be up to 10 times faster. 4) The performance decline over time, particularly noticeable in the CADE method, suggests that continuous adaptation might be necessary not just for classifiers, but also for detection models.
4.2 Challenges in Human-in-the-loop
@@
In the active learning framework utilized for drift adaptation, the involvement of human annotation presents specific challenges. The primary concern is the necessity to maintain a low human labeling budget, crucial due to the extensive efforts required in malware analysis [43]. The current adaptation approach, predominantly relying on retraining, is far from ideal in this context. This situation calls for a more refined technique that optimally utilizes human input while minimizing its extent, thereby achieving a balance between expert knowledge and resource constraints. Additionally, the potential for errors in human labeling, similar to inaccuracies observed in automated tools like antivirus scanners [71], underscores the need for resilience in the adaptation technique to mitigate labeling inconsistencies.
@@ 5.1 Design Insights
5.2 Concept-based Drift Detection
Model Sensitive Concept Learning. We explore the idea of concept learning to enhance contrastive learning for latent representations in the drift detector. In the context of machine learning, concept learning typically involves the task of inferring boolean-valued functions from labeled examples to generalize specific instances to broader concepts [31, 16]. It is a form of supervised learning, where each concept, such as distinct geometric shapes (e.g., circles, squares, and triangles), is associated with a binary label for tasks like object identification [9]. Our intuition is that learning generalized concepts can offer a more direct method for identifying concept drift. To this end, we first define the concepts as malicious behaviors (e.g., information stealing, remote control, and stealthy download) under the malware context. Subsequently, our goal is to integrate the supervised concept learning with the unsupervised contrastive learning.
@@ 5.2 Concept-based Drift Detection
Concept Reliability Detector. With the trained detector, we can leverage it to detect drifting samples for incoming test data. Our design draws inspiration from a notable observation in the state-of-the-art HCC detector. Considering the significant contribution of the cross-entropy-based pseudo loss to the detection performance, we focus on enhancing the contrastive-based element in Equation 7. We firstly achieve this by maintaining the neighborhood function while substituting ^$\hat{\mathcal{L}}_{hc}$ with ^$\hat{\mathcal{L}}_{ce}$, leading to the introduction of the neighborhood cross-entropy (NCE) based pseudo loss, defined as
@@
In line with the experimental setup described in Section 4.1, we observe that this adjustment can improve the HCC detector’s performance to some extent: as shown in Section 5.2, the improvement in the average AUC achieves $6.11\%$.
@@
Table 3: The detection AUC of the proposed NCE compared with the original ce-based detector in HCC.
To comprehend the underlying principle, it’s important to understand that NCE calculates uncertainty for nearest neighbors based on the assumption that a non-drifting sample should remain certain even if slightly perturbed. This leads us to the samples reconstructed by the autoencoder, which essentially represent another type of perturbation. Consequently, the pseudo loss defined with our concept reliability loss (Equation 9) can align well with this principle. The drift scoring function of our detector is
@@
5.3 Explanatory Drift Adaptation
@@
We enable human analysts to provide feedback not only on the predicted labels of drifting samples but also on their behavioral explanations. For instance, in malware classification tasks, analysts might encounter a drifting sample identified as belonging to the GhostCtrl family, exhibiting behaviors like PrivacyStealing, SMSCALL, RemoteControl, and Ransom. Feedback in this scenario could be formatted as: {mdframed}[style=mystyle]
@@
The feedback can be generalized to identify specific malware families with minimal annotation effort, as in our experiments. An example of this is presented in Appendix C.
@@ 6 System Evaluation
6.1 Experimental Setup
Dataset. We employ two malware datasets for malware family classification tasks, as illustrated in Table 10. Our first dataset is the well-known Drebin dataset [4], adhering to the same criteria used in the CADE paper. To capture more recent trends in the evolving malware landscape, we utilize the Malradar dataset [69]. For each dataset, we select $8$ families, each comprising at least $100$ malware samples. The resulting datasets consist of $3,317$ and $2,589$ malware samples for Drebin and Malradar, respectively. Both datasets are well-labeled with respect to malware families. Additionally, Malradar provides behavioral labels derived from threat reports for each family, and we select $10$ distinct behaviors to form its behavioral labels. To ensure consistency in behavioral analysis, we augment the Drebin dataset with these same behaviors. Despite the simplicity, this approach is in line with an active learning setting that typically utilizes limited human effort. For more details about the two datasets, please also refer to Appendix B.
@@ 6.1 Experimental Setup
Table 4: Drift adaptation results on two malware datasets within the three features.
@@
Table 5: Drift adaptation under label noises with budget 50.
6.2 Drift Detection
Baseline and Metric. For assessing our drift detector’s performance in inter-class scenarios, we adopt the baseline methods previously utilized in the intra-drift detection, as outlined in Section 4.1. However, the HCC detector is excluded from this evaluation, considering its incompatibility with constructing hierarchical contrastive loss in these contexts. To ensure a fair comparison, both the CADE detector and our detector, which each leverage an autoencoder model, are configured to share the same architecture across all features (detailed in Appendix D). For the metric, we utilize the AUC calculated from the detector’s drifting score output and the ground truth labels, which are determined by the held-out malware families during the training process.

### v2

# Dream: Combating Concept Drift with Explanatory Detection and Adaptation in Malware Classification

@@ Abstract
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2025 https://dx.doi.org/10.14722/ndss.2025.xxxxx www.ndss-symposium.org
I Introduction
@@
Existing methods for drift detection, however, fall short in two main aspects. Firstly, they often falsely neglect the patterns which the targeted classifier depends on. For example, the CADE detector [7] leverages an independent autoencoder to learn a distance function, identifying drifts by the distances with training data. Such misalignment can lead to inefficiency in detecting model-specific drifts, especially when dealing with complex classifier feature spaces, as validated by our experiments in Section VI-B. Secondly, a constant involvement of training data during the testing phase introduces practical challenges. For instance, methods use training points to query reference uncertainties [6, 5] or search for nearest neighbors [8]. Such practices are problematic particularly in local deployment scenarios, where managing large training datasets raises storage and security concerns [10].
@@
We develop a system, called Dream, distinguished by concept-based drift detection and explanatory adaption for malware classification. To accurately learn malware concepts, we combine supervised concept learning [15] with unsupervised contrastive learning [16]. Our drift detector is thus a semi-supervised model structured with an autoencoder. Within this model, elements in the latent space are representative of behavioral concepts that can be labeled, such as remote control and stealthy download. Specifically, Dream achieves model sensitive and training data independent drift detection with the concept reliability measure. This unique measure is based on the intuition that if a sample is not subject to drift, then the concepts embedded in our detector should be reliable; consistency should be reflected in similar classifier predictions for the original and reconstructed samples.
@@
We evaluate Dream with two distinct malware datasets [7, 19] and three state-of-the-art classifiers [20, 21, 22]. Evaluation results demonstrate that Dream outperforms three existing drift detection methods [6, 7, 8] in malware classification, achieving an average improvement of 11.5%, 12.0%, and 13.6% in terms of the AUC  [^1] The AUC specifically refers to the Area Under the Receiver Operating Characteristic (ROC) Curve [23] in this paper. metric. For drift adaptation, Dream notably enhances the common retraining-based approach, increasing the F1-score by 168.4%, 73.5%, 22.2%, 19.7%, and 6.0% when the human analysis budget is set at $10$, $20$, $30$, $40$, and $100$, respectively. A standout observation is our increased advantage at lower human labeling costs. For instance, analysts can analyze over 80% fewer samples while still achieving a test accuracy of $0.9$ for the updated classifier.
@@
II Malware Concept Drift
In this section, we delve into the background of Android malware concept drift. Firstly, we discuss two types of drift for learning-based malware applications. Then, we outline the widely employed techniques for drift detection and drift adaptation (with supplements available in Section IX). Finally, we highlight key studies in this domain, underscoring their unique research orientations and inherent limitations.
II-A Root Cause and Categorization
@@
TABLE I: Summary of existing work in terms of the detection method, adaptation method, and the applicable scope.
II-B Drift Detection and Adaptation
Drift detection is essential for stakeholders to determine when to update the classifier. Two primary Out-of-Distribution (OOD) detection methods [31] are chiefly employed. 1) Uncertainty Estimation gauges a model’s prediction confidence. Techniques like Bayesian Neural Networks [32] and Deep Ensembles [33] inherently quantify this measure, but they are specific to model architectures. Standard DNNs often use heuristics like softmax-derived uncertainty, which can be misleading as overfit DNNs may assign high softmax values to unfamiliar data, falsely indicating confidence [34, 35]. 2) Nonconformity Scoring, in contrast, evaluates how unusual new data is compared to a calibration set [5]. It involves a distance function and a predictor for representing calibration and test data. While traditional techniques depend heavily on the calibration set, recent approaches use contrastive learning to refine distance functions [7, 8, 36]. However, their model independent nature can lead to inconsistent results across different architectures.
Drift adaptation in malware classification traditionally involves redesigning feature spaces, which would require expensive analysis and training efforts [37, 38]. To satisfy online scenarios, some researches adopt a conservative approach, named classification with rejection [5, 6]. In this approach, ambiguous samples are temporarily withheld for expert analysis to minimize misclassification risks. While providing a safeguard against immediate threats, its efficacy in long-term resilience is limited. Recent strategies, inspired by active learning, introduce an analysis budget to prioritize a subset of drifting samples for human review and subsequent integration into retraining [8]. However, it’s notable that existing studies all focus on the adaptation for intra-class drift, where the number of output classes is unchanged. Additionally, there’s growing interest in drift explanation, where the focus is on understanding the reasons behind drift samples to streamline manual analysis [7, 39, 40]. Nevertheless, the post-analysis application of these explanations in updating classifiers remains underexplored.
II-C Summary of Existing Work
Advancements in drift detection and adaptation techniques are crucial for combating malware drifts effectively. Focusing on these aspects, three notable studies have contributed significantly to the field. As outlined in Table I, we discuss their methodologies and comparative research focuses below, leaving technical details of their detectors in Section IV-A.
@@ II-C Summary of Existing Work
III Problem Setup
In this section, we define our research scope and formalize three important components in the active learning-based concept drift mitigation strategy.
III-A Research Scope
@@
Distinguishing our method from HCC, we emphasize model-agnostic techniques, which are adept at operating effectively without modifications to the underlying classifier’s architecture. Recognizing that intra-class drift scenarios have already been extensively explored, our work primarily addresses the challenges associated with inter-class drifts. That is, we aim to delve deeply into the detection and adaptation strategies for drifts caused by unseen malware families. Nevertheless, it’s important to note that our proposed solution is not exclusively tailored to addressing inter-class drifts. The generalizability of our approach will also undergo a detailed evaluation in Section VII.
@@ III-B Formalization
IV Motivation and Challenge
IV-A Characterizing Current Drift Detectors
@@
TABLE II: Intra-class drift detection performance of existing works. For the state-of-the-art method HCC, we also make a separate study of the two items in its drift scoring function.
IV-B Lessons from Current Detectors
@@
Evaluation Method. For malware detection tasks, we leverage malware samples from the Malradar dataset [19], which covers 4,410 malware across 180 families including singletons. Moreover, we make efforts to collect 43,641 benign samples in the same year period from 2015 to 2020. Our dataset has been carefully designed to mitigate spatial bias and temporal bias [1]. For the classifier, we use the Drebin features and an MLP model (detailed in Section VI-A). Regarding baselines, we take into account the vanilla probability-based detector (as depicted in Equation 4) and the three innovative detectors formalized in Section IV-A. During the evaluation phase, we assign positive drift labels to samples where predictions are incorrect, following the rationale that these misclassified samples are particularly valuable for adapting the classifier. This approach differs from inter-class drift detection, where positive drift labels are typically assigned based on the presence of unseen families, as utilized in existing work [7]. Then we compute the AUC metric using drifting scores and the ground truth drift labels.
@@
Key Findings. Our evaluation results are shown in Table II, from which we have identified four key findings. 1) The performance order in our experiments, i.e., HCC, Transcendent, and Probability all outperforming CADE, highlights the essential role of model sensitivity in effective drift detection. 2) Although CADE is not specifically tailored for intra-class drift, its superior performance in a non-model sensitive comparison with the $\text{HCC}_{hc}$ indicates the potential advantages of data autonomy and its autoencoder-based structure. 3) HCC surpassing Transcendent in performance indicates that learning a distance metric, rather than relying on calibration, can be more effective and efficient in complex models like DNNs. This observation contrasts with Transcendent’s findings on SVM models and is further supported by the efficiency of the detector’s training process, which can be up to 10 times faster. 4) The performance decline over time, particularly noticeable in the CADE method, suggests that continuous adaptation might be necessary not just for classifiers, but also for detection models.
IV-C Challenges in Human-in-the-loop
@@
In the active learning framework utilized for drift adaptation, the involvement of human annotation presents specific challenges. The primary concern is the necessity to maintain a low human labeling budget (i.e., the number of newly labeled samples to achieve acceptable accuracy), crucial due to the extensive efforts required in malware analysis [11]. The current adaptation approach, predominantly relying on retraining, is far from ideal in this context. This situation calls for a more refined technique that optimally utilizes human input while minimizing its extent, thereby achieving a balance between expert knowledge and resource constraints. Additionally, the potential for errors in human labeling, similar to inaccuracies observed in automated tools like antivirus scanners [13], underscores the need for resilience in the adaptation technique to mitigate labeling inconsistencies.
@@ V-A Design Insights
V-B Concept-based Drift Detection
Model Sensitive Concept Learning. We explore the idea of concept learning to enhance contrastive learning for latent representations in the drift detector. In the context of machine learning, concept learning typically involves the task of inferring boolean-valued functions from labeled examples to generalize specific instances to broader concepts [15, 47]. Our intuition is that learning generalized concepts can offer a more direct method for identifying concept drift. To this end, as illustrated in Figure 3, we first define the concepts as malicious behaviors (e.g., information stealing, remote control, and stealthy download) under the malware context. Subsequently, our goal is to integrate the supervised concept learning with the unsupervised contrastive learning.
Figure 3: Model sensitive concept learning. To enhance contrastive autoencoder-based drift detector with model sensitivity, Dream enriches the latent space with concepts of malicious behaviors, and it learns those concepts with presence and reliability based loss items.
@@ V-B Concept-based Drift Detection
Data Autonomous Detector. With the trained detector, we can leverage it to detect drifting samples for incoming test data. Our design draws inspiration from a notable observation in the state-of-the-art HCC detector. Considering the significant contribution of the cross-entropy-based pseudo loss to the detection performance, we focus on enhancing the contrastive-based element in Equation 7. We firstly achieve this by maintaining the neighborhood function while substituting ^$\hat{\mathcal{L}}_{hc}$ with ^$\hat{\mathcal{L}}_{ce}$, leading to the introduction of the neighborhood cross-entropy (NCE) based pseudo loss, defined as
@@
In line with the experimental setup described in Section IV-B, we observe that this adjustment can improve the HCC detector’s performance to some extent: as shown in Section V-B, the improvement in the average AUC achieves $6.11\%$.
@@
TABLE III: The detection AUC of the proposed NCE compared with the original ce-based detector in HCC.
To comprehend the underlying principle, it’s important to understand that NCE calculates uncertainty for nearest neighbors based on the assumption that a non-drifting sample should remain certain even if slightly perturbed. This leads us to the samples reconstructed by the autoencoder, which essentially represent another type of perturbation. Consequently, the pseudo loss defined with our concept reliability loss (Equation 9) can align well with this principle. The drift scoring function of our Concept Reliability Detector is
@@
V-C Explanatory Drift Adaptation
As shown in Figure 4, Dream addresses the human-in-the-loop challenges with 1) intelligible explanations that can facilitate the behavior labeling process to operate on the classifier, and 2) efficient drift adaption algorithm that can decrease the labeling budget required for the accuracy of updated classifier. We discuss the two solutions below.
Figure 4: Human-in-the-loop solutions in Dream.
@@ V-C Explanatory Drift Adaptation
We enable human analysts to provide feedback not only on the predicted labels of drifting samples but also on their behavioral explanations. For instance, in malware classification tasks, analysts might encounter a drifting sample identified as belonging to the GhostCtrl family, exhibiting malicious behaviors like PrivacyStealing, SMSCALL, RemoteControl, and Ransom. Feedback in this scenario can be formatted as: {mdframed}[style=mystyle]
@@
The feedback can be generalized to identify specific malware families with minimal annotation effort, as in our experiments. An example of this is presented in Section -C.
@@ VI System Evaluation
VI-A Experimental Setup
Dataset. We employ two malware datasets for malware family classification tasks, as illustrated in Table X. Our first dataset is the well-known Drebin dataset [20], adhering to the same criteria used in the CADE paper. To capture more recent trends in the evolving malware landscape, we utilize the Malradar dataset [19]. For each dataset, we select $8$ families, each comprising at least $100$ malware samples. The resulting datasets consist of $3,317$ and $2,589$ malware samples for Drebin and Malradar, respectively. Both datasets are well-labeled with respect to malware families. Additionally, Malradar provides behavioral labels derived from threat reports for each family, and we select $10$ distinct behaviors to form its behavioral labels. To ensure consistency in behavioral analysis, we augment the Drebin dataset with these same behaviors. Despite the simplicity, this approach is in line with an active learning setting that typically utilizes limited human effort. For more details about the two datasets, please also refer to Section -B.
@@ VI-A Experimental Setup
(a) Transcendent, Drebin
(b) CADE, Drebin
(c) Probability, Drebin
(d) (Ours) Dream, Drebin
(e) Transcendent, Mamadroid
(f) CADE, Mamadroid
(g) Probability, Mamadroid
(h) (Ours) Dream, Mamadroid
(i) Transcendent, Damd
(j) CADE, Damd
(k) Probability, Damd
(l) (Ours) Dream, Damd
@@
TABLE IV: Drift adaptation results on two malware datasets within the three features.
@@
TABLE V: Drift adaptation under label noises with budget 50.
VI-B Drift Detection
Baseline and Metric. For assessing our drift detector’s performance in inter-class scenarios, we adopt the baseline methods previously utilized in the intra-drift detection, as outlined in Section IV-B. However, the HCC detector is excluded from this evaluation, considering its incompatibility with constructing hierarchical contrastive loss in these contexts. To ensure a fair comparison, both the CADE detector and our detector, which each leverage an autoencoder model, are configured to share the same architecture across all features (detailed in Section -D). For the metric, we utilize the AUC calculated from the detector’s drifting score output and the ground truth labels, which are determined by the held-out malware families during the training process.